OpenStudy (anonymous):

Divide. Move the decimal points when necessary. 12.456/1.8

OpenStudy (anonymous):

\[\frac{12.456}{1.8}=\frac{12.456*10}{1.8*10}=\frac{124.56}{18}=\frac{124.56*100}{18*100}=\frac{12456}{1800}=\frac{173}{25}\]